# SelaVPR++: Towards Seamless Adaptation of Foundation Models for Efficient Place Recognition

### v1

@@ Abstract
Index Terms:
Visual place recognition, foundation models, parameter-efficient transfer learning, deep hashing.
@@ I Introduction
Figure 1: Attention map visualizations of the pre-trained foundation model (DINOv2), SelaVPR, and SelaVPR++. The pre-trained model pays attention to some regions that are useless to identify places, e.g., dynamic riders. SelaVPR and SelaVPR++ focus on discriminative regions (buildings and trees). Compared with SelaVPR, SelaVPR++ focuses on more landmarks (e.g., smokestack) and eliminates more dynamic interference (e.g., truck), i.e., performs better in details.
@@
Figure 2: Comparison between different transfer learning methods. (a) is the common full fine-tuning, in which all blocks are trainable. (b) is a popular PETL method, where only inner adapters are trainable. But the backpropagation still passes through the entire frozen backbone. (c) is our memory-efficient adaptation, which reduces training memory usage by eliminating the need of backpropagation through the backbone. For the sake of clarity, we do not show the residual connection parallel to the adapter in (c).
@@
Nevertheless, SelaVPR still can be further improved in terms of efficiency and performance. First, although adding the tunable lightweight adapters to the frozen pre-trained model achieves parameter-efficient adaptation, it is not efficient in training time and GPU memory usage. This is because the gradient computation for trainable parameters still requires backpropagation through the frozen backbone, as shown in Fig. 2 (b), which results in substantial demands on training time and GPU memory. Second, the re-ranking based on dense local features typically incurs huge overhead in time, memory, and storage. Although SelaVPR significantly reduces the retrieval latency compared to common two-stage methods by eliminating the geometric verification requirement, it still requires huge storage to save local features. This limits its application in resource-constrained and large-scale VPR scenes. Additionally, with the development of one-stage VPR methods [23, 24, 25, 26], especially training VPR models on large-scale datasets with full supervision [16, 26] (instead of weakly supervised training as in SelaVPR), using only global features can also achieve relatively promising results, implying that the paradigm of using local features for re-ranking, which consumes significant resources but offers limited performance gains, is not economical.
@@
To address these issues, we further improve our ICLR 2024 work SelaVPR, and demonstrate a novel insight towards seamless adaptation of foundation models for the more effective and efficient VPR model designs, named SelaVPR++. To begin with, instead of inserting the adapters within the backbone, we connect them to the backbone in parallel and use them to progressively refine the intermediate features from the transformer blocks of the frozen backbone, as shown in Fig. 2 (c). This can avoid the gradient backpropagation through the frozen backbone, thus significantly reducing computation and lowering GPU memory footprint during training. In addition, to incorporate local priors beneficial for VPR and make up for the shortcoming of the vanilla adapter, which operates solely on the channel dimension of intermediate features without modeling interactions along the spatial axes, we introduce the multi-scale convolution (MultiConv) adapter to upgrade the adapter used in SelaVPR. Then, we propose an innovative two-stage paradigm for more efficient VPR. Rather than relying on local features for time- and memory-consuming re-ranking as existing methods do, we employ compact low-dimensional (e.g., 512-dim) binary features for initial retrieval to get top-k candidate images and robust high-dimensional (e.g., 4096-dim) floating-point features for re-ranking. That is, our method employs global features for both stages, without any local features. Specifically, we utilize deep hashing to obtain compact binary features, in which we design a similarity-constrained quantization loss that only preserves the similarity consistency of feature pairs before and after quantization to address the conflict between common quantization loss and metric loss. We also combine it with straight-through estimation [27], which can tackle the gradient issue of quantization operation, for better performance. Since our method computes the Hamming distance between binary features to provide candidates, it can achieve faster retrieval speed than current one-stage methods. Finally, we improve the training strategy and unify the training protocol of several commonly used VPR training datasets. Unlike training on individual datasets with weak supervision in SelaVPR, we combine several datasets in the unified protocol for fully supervised training. Our SelaVPR++ work brings the following main contributions (to the conference version SelaVPR):
@@
1) We propose a novel method to achieve seamless and efficient adaptation of foundation models for VPR, in which trainable lightweight MultiConv adapters are used to refine intermediate features from the frozen foundation model (without the gradient backpropagation through it in training), thereby realizing parameter-, time- and memory-efficient adaptation. Meanwhile, the MultiConv adapter not only models spatial interactions between patch tokens but also introduces multi-scale local priors to the VPR model.
@@ II-A One-Stage VPR
The most common VPR approach performs nearest neighbor search using global features to find the most similar place images without considering re-ranking, which is also recognized as one-stage (i.e., global retrieval) VPR. In the early VPR methods, the global features were commonly produced using aggregation algorithms, such as Bag of Words [30] and VLAD [8], to process the traditional hand-crafted features (e.g., SIFT [31], SURF [32]). With the advancement of deep learning techniques, many works [33, 34, 35, 20, 36, 37, 38, 39, 40, 41, 42, 43, 44, 23] have employed various deep features for the VPR task. Some works integrated the aggregation methods into neural networks [6, 45, 46], and improved training strategies [47, 28, 16], to achieve better performance. Nevertheless, most one-stage VPR approaches are susceptible to perceptual aliasing due to the use of aggregated features while neglecting spatial information. One recent work [48] first used pre-trained foundation models for the VPR task. However, this work did not perform any fine-tuning, making it difficult to fully unleash the capability of these models for VPR. Some follow-up works [24, 26] attempted to directly fine-tune (the last multiple blocks of) foundation models, but this way is primarily suitable for medium-size models, where performance may decrease as the model size increases, i.e. hard to train large models [24].
@@ II-B Two-Stage VPR
The two-stage (i.e., hierarchical) VPR methods with re-ranking [49, 50, 11, 51, 52, 12, 53, 54, 55, 13, 56] have been proven as an effective way to further improve performance. These approaches typically retrieved top-k candidate images over the whole database using compact global feature representation, such as NetVLAD [6] or Generalized Mean (GeM) pooling [57], then re-ranked candidates by performing local matching between the query image and each candidate using local descriptors. However, most of these methods required geometric consistency verification after local matching [7, 11, 12, 56] or taking into account spatial constraints during matching [51, 53, 55], which greatly increases the computational latency. In our SelaVPR [13] work, we fine-tuned a foundation model with a local adaptation module to obtain dense local features, which can be directly used in cross-matching for re-ranking, without time-consuming geometric verification. However, SelaVPR still required substantial storage (and memory) for local features in re-ranking as previous methods, restricting its applicability in resource-constrained and large-scale VPR scenarios. Additionally, with the rapid development of one-stage VPR [23, 24, 26], spending substantial time and resources on re-ranking with local features for very limited performance gains has become inefficient and uneconomical. Therefore, in this work, we attempt to develop an innovative two-stage VPR paradigm that performs initial retrieval using compact low-dimensional binary features derived from deep hashing, and then utilizes high-dimensional floating-point features for re-ranking, significantly reducing time, memory, and storage consumption. Although a few previous works [58, 59] have used deep hashing in VPR, they used the sigmoid function to approximate the binarization during training, which is hard to train and has been replaced by the method that imposes the quantization loss (a regularizer) [60]. More importantly, they used binary features for direct retrieval instead of candidate retrieval with re-ranking, which did not achieve good performance. To our best knowledge, we are the first to apply binary descriptors for initial retrieval in two-stage VPR.
@@ II-C Parameter-Efficient Transfer Learning
Recent work [17, 61, 62, 19] demonstrated the visual foundation model can produce powerful feature representation and achieve excellent performance on multiple tasks. These works commonly trained the ViT [63] model or its variants with large quantities of parameters on huge amounts of data. The parameter-efficient transfer learning (PETL) [21], a.k.a. parameter-efficient fine-tuning (PEFT), first proposed in natural language processing, is an effective way to adapt foundation models to various downstream tasks, which can reduce trainable parameters and avoid catastrophic forgetting. The main PETL methods fall broadly into three categories: adding task-specific adapters [21], prompt tuning [22], and Low-Rank Adaptation (LoRA) [64]. Our previous works [13, 23] followed the first to adapt the pre-trained foundation models to the VPR task, while they are efficient only in terms of parameters, not in training time and GPU memory usage. Fortunately, there are some recent works [65, 66] that have preliminarily attempted to explore memory-efficient transfer learning methods. Inspired by them, we design a memory-efficient adaptation architecture tailored for the VPR task. Concretely, we use the tunable adapters to refine the intermediate features from the frozen backbone instead of inserting them into it to avoid backpropagation through the backbone, and upgrade vanilla adapters to MultiConv adapters to introduce the local priors and facilitate feature interactions along the spatial axes. As a result, we can achieve parameter-, time-, and memory-efficient adaptation and get highly robust VPR models.
@@ III-A Preliminary
where $x_{l-1}$ and $x_{l}$ are the output of the $(l-1)$-th and $l$-th transformer block.
For the feature map output by CNN models, a common practice of the VPR method is to use NetVLAD [6] or GeM pooling [57] to aggregate it into a global feature to conduct nearest neighbor search (i.e., global retrieval). For the ViT model, the output consists of one class token and $N$ patch tokens, where the class token can be directly used as the global feature to represent places. Meanwhile, $N$ patch tokens can also be reshaped as a feature map (similar to CNN). In this work, instead of using the class token, we transform the feature map into the global feature using GeM pooling.
@@ III-B Memory-Efficient MultiConv Adaptation
However, due to the adapters inserted in the transformer block, this will lead to significant GPU memory consumption during training. Here, we conduct a brief equation derivation and theoretical analysis. Consider a transformer network with $L$ blocks, each containing the frozen parameters $\theta_{i}$ (frozen modules) and trainable parameters $\alpha_{i}$ (trainable modules), and producing the output $x_{i}$, which depend on both parameters $\theta_{i},\alpha_{i}$. What we need to do is to minimize a loss function $\mathcal{L}$ based on stochastic gradient descent. To be specific, the gradient for the trainable parameters $\alpha_{i}$ is calculated using the chain rule during backpropagation, formulated as follows:
@@
To this end, inspired by previous works [65, 66], we introduce a distinctive adaptation method. Instead of inserting the adapters into the backbone, we use a parallel network that does not require gradient backpropagation through the backbone during training as shown in Fig. 2 (c). It is a simple yet effective architecture, which directly takes the intermediate features from the frozen backbone as input and progressively refines them for better place representation. Concretely, we adapt the ViT-based foundation model DINOv2 [19] as the backbone, which is composed of a patch embedding layer and $L$ transformer blocks, and thereby $L+1$ intermediate features, $x_{0},x_{1},x_{2}...,x_{L}$, each consisting of $N+1$ tokens with a dimension of $D$, so $x_{i}\in\mathcal{R}^{(N+1)\times D}$ as mentioned in III-A. To refine them progressively, we construct the parallel adapters alongside the backbone. Meanwhile, we add residual connections to each adapter for better performance. We represent the output of the $l$-th parallel adapter as $y_{l}$. The whole computation can be formulated as:
@@
In this work, we introduce the multi-scale convolution (MultiConv) adapter, which is first proposed in our previous CricaVPR work [23] but is used by inserting it into the transformer block (similar to SelaVPR but different from SelaVPR++). Notably, in CricaVPR, the MultiConv adapter is used to add multi-scale local prior knowledge to the model for enhancing the performance in the VPR task. However, in SelaVPR++, it not only introduces multi-scale local priors, but more importantly, it breaks the limitation that the vanilla adapter cannot model the interactions between different patch tokens. In the MultiConv adapter, we add a MultiConv module (with a skip connection) between the activation layer and the up-projection layer, as shown in Fig. 3 (c). This module contains three parallel convolution paths of different scales (1$\times$1, 3$\times$3, 5$\times$5), which is inspired by the inception module in GoogLeNet [69]. The 1$\times$1 convolution is also applied before the 3$\times$3 and 5$\times$5 convolutions to reduce the number of channels. This design and the bottleneck structure make our MultiConv adapter still lightweight. The outputs of the three convolution paths are subsequently concatenated to achieve integration. The MultiConv adapter only processes patch tokens, which are reshaped into a $W\times H\times D$-dim (from $N\times D$) feature map to restore the spatial position before being input into the adapter. In this way, the spatially adjacent patch tokens can be modeled interactions via 3$\times$3 and 5$\times$5 convolutions. Moreover, since the MultiConv adapter has introduced multi-scale local prior knowledge to the model, we no longer need to purposefully design a separate local adaptation module to process local information for VPR as in our previous SelaVPR.
@@ III-D Similarity-Constrained Deep Hashing
Hashing has gained widespread attention in the approximate nearest neighbor search for large-scale image retrieval due to its computational efficiency and retrieval quality by encoding data into binary vectors and performing search with Hamming distance. Deep hashing, in particular, has further enhanced performance by combining deep representation learning with hash coding. In this work, we attempt to adopt deep hashing to provide compact binary place representation, which is achieved by using a sign function $sgn(\cdot)$ to quantize floating-point features $\boldsymbol{f}_{i}$} into binary hash codes $\boldsymbol{b}_{i}$}. Formally,
@@
In summary, to train a deep hashing model, we can use the total loss $\mathcal{L}$ that combines the metric loss $\mathcal{L}_{M}(\boldsymbol{f})$ and similarity-constrained loss $\mathcal{L}_{Q}(\boldsymbol{f},\boldsymbol{b})$ by a weight $\lambda$ as:
@@
Another solution to address the gradient issue of quantization operation is straight-through estimation (STE) [27]. We employ it to perform the backpropagation of deep hashing in the VPR task. Specifically, we use the original sign function $sgn(\cdot)$ during forward propagation to obtain the binary hash codes. In back propagation, the derivative of binary hashing operation is replaced with the derivative of a smooth approximation $g$, which can be formulated as:

### v2

@@ I Introduction
Figure 1: Heatmap visualizations of feature maps from the pre-trained foundation model, SelaVPR, and SelaVPR++. The pre-trained model pays attention to some regions that are useless for VPR, e.g., dynamic riders. SelaVPR and SelaVPR++ focus on discriminative regions (buildings and trees). Compared with SelaVPR, SelaVPR++ focuses on more landmarks (e.g., smokestack) and eliminates more dynamic interference (e.g., truck), i.e., performs better in detail.
@@
Figure 2: Comparison between different transfer learning methods. (a) is the common full fine-tuning, in which all blocks are trainable. (b) is a popular PETL method, where only inner adapters are trainable. But the backpropagation still passes through the entire frozen backbone. (c) is the memory-efficient adaptation following the basic framework of previous work [23], which reduces training memory usage by eliminating the need for backpropagation through the backbone.
@@
Nevertheless, SelaVPR can still be further improved in terms of efficiency and performance. First, although adding the tunable lightweight adapters to the frozen pre-trained model achieves parameter-efficient adaptation, it is not efficient in training time and GPU memory usage. This is because the gradient computation for trainable parameters still requires backpropagation through the frozen backbone, as shown in Fig. 2 (b), which results in substantial demands on training time and GPU memory. Second, the re-ranking based on dense local features typically incurs huge overhead in time, memory, and storage. Although SelaVPR significantly reduces the retrieval latency compared to common two-stage methods by eliminating the geometric verification requirement, it still requires huge storage to save local features. This limits its application in resource-constrained and large-scale VPR scenes. Additionally, with the development of one-stage VPR methods [24, 25, 26, 27], especially training VPR models on large-scale datasets with full supervision [16, 27] (instead of weakly supervised training as in SelaVPR), using only global features can also achieve relatively promising results, implying that the paradigm of using local features for re-ranking, which consumes significant resources but offers limited performance gains, is not economical.
@@
To address these issues, we further improve our SelaVPR, and demonstrate a novel insight towards seamless adaptation of foundation models for the more effective and efficient VPR model designs, named SelaVPR++. To begin with, instead of inserting the adapters within the backbone, we connect them to the backbone in parallel and use them to progressively refine the intermediate features from the transformer blocks of the frozen backbone following the basic framework in the work [23], as shown in Fig. 2 (c). This can avoid the gradient backpropagation through the frozen backbone, thus significantly reducing computation and lowering GPU memory footprint during training. In addition, to incorporate local priors beneficial for VPR and make up for the shortcoming of the vanilla adapter, which operates solely on the channel dimension of intermediate features without modeling interactions along the spatial axes, we introduce the multi-scale convolution (MultiConv) adapter to upgrade the adapter used in SelaVPR. Then, we propose an innovative two-stage paradigm for more efficient VPR. Rather than relying on local features for time- and memory-consuming re-ranking as existing methods do, we use compact low-dimensional (e.g., 512-dim) binary features for initial retrieval to get top-k candidate images and robust high-dimensional (e.g., 4096-dim) floating-point features for re-ranking. That is, our method employs global features for both stages, without any local features. Specifically, we utilize deep hashing to obtain compact binary features, in which we design a similarity-constrained quantization loss that only preserves the similarity consistency of feature pairs before and after quantization to address the conflict between common quantization loss and metric loss. We also combine it with straight-through estimation [28], which can tackle the gradient issue of quantization operation, for better performance. Since our method computes the Hamming distance between binary features to provide candidates, it can achieve faster retrieval speed than current one-stage methods. Finally, we improve the training strategy and unify the training protocol of several commonly used VPR training datasets. Unlike training on individual datasets with weak supervision in SelaVPR, we combine several datasets in the unified protocol for fully supervised training. Our SelaVPR++ work brings the following main contributions (to the previous SelaVPR):
@@
1) Inspired by the previous memory-efficient adaptation paradigm [23, 29], we propose a novel method to achieve seamless and efficient adaptation of foundation models for VPR, in which trainable lightweight MultiConv adapters are used to refine intermediate features from the frozen foundation model (without the gradient backpropagation through it in training), thereby realizing parameter-, time- and memory-efficient adaptation. Meanwhile, the MultiConv adapter not only models spatial interactions between patch tokens but also introduces multi-scale local priors to the VPR model.
@@ II-A One-Stage VPR
The most common VPR approach performs nearest neighbor search using global features to find the most similar place images without considering re-ranking, which is also recognized as one-stage (i.e., global retrieval) VPR. In the early VPR methods, the global features were commonly produced using aggregation algorithms, such as Bag of Words [32] and VLAD [8], to process the traditional hand-crafted features (e.g., SIFT [33], SURF [34]). With the advancement of deep learning techniques, many works [35, 36, 37, 20, 38, 39, 40, 41, 42, 43, 44, 45, 46, 24] have employed various deep features for the VPR task. For instance, some works integrated the aggregation methods into neural networks [6, 47, 48] and improved training strategies [49, 30, 16] to achieve better performance, and other studies [50, 51] explored model quantization to reduce memory consumption and latency during deep feature extraction. Nevertheless, most one-stage VPR approaches are susceptible to perceptual aliasing due to the use of aggregated features while neglecting spatial information. One recent work [52] first used pre-trained foundation models for the VPR task. However, this work did not perform any fine-tuning, making it difficult to fully unleash the capability of these models for VPR. Some follow-up works [25, 27] attempted to directly fine-tune (the last multiple blocks of) foundation models, but this way is primarily suitable for medium-size models, where performance may decrease as the model size increases, i.e. hard to train large models [25].
@@ II-B Two-Stage VPR
The two-stage (i.e., hierarchical) VPR methods with re-ranking [53, 54, 11, 55, 56, 12, 57, 58, 59, 13, 60] have been proven as an effective way to further improve performance. These approaches typically retrieved top-k candidate images over the whole database using compact global feature representation, such as NetVLAD [6] or Generalized Mean (GeM) pooling [61], then re-ranked candidates by performing local matching between the query image and each candidate using local descriptors. However, most of these methods required geometric consistency verification after local matching [7, 11, 12, 60] or taking into account spatial constraints during matching [55, 57, 59], which greatly increases the computational latency. In our SelaVPR [13] work, we fine-tuned a foundation model with a local adaptation module to obtain dense local features, which can be directly used in cross-matching for re-ranking, without time-consuming geometric verification. However, SelaVPR still required substantial storage (and memory) for local features in re-ranking as previous methods, restricting its applicability in resource-constrained and large-scale VPR scenarios. Additionally, with the rapid development of one-stage VPR [24, 25, 27], spending substantial time and resources on re-ranking with local features for very limited performance gains has become inefficient and uneconomical. So, in this work, we attempt to develop an innovative two-stage VPR paradigm that performs initial retrieval using compact low-dimensional binary features derived from deep hashing, and then utilizes high-dimensional floating-point features for re-ranking, significantly reducing time, memory, and storage consumption. Although a few previous works [62, 63, 64, 65] have used hashing or deep hashing in VPR and similar tasks, they were either not end-to-end trained, or used the sigmoid function to approximate the binarization during training, which is hard to train and has been replaced by the method that imposes the quantization loss (a regularizer) [66]. More importantly, they used binary features for direct retrieval instead of candidate retrieval with re-ranking, which did not achieve good performance. To our best knowledge, we are the first to apply binary descriptors for initial retrieval in two-stage VPR.
@@ II-C Parameter-Efficient Transfer Learning
Recent work [17, 67, 68, 19] demonstrated that visual foundation models can produce powerful feature representations and achieve excellent performance on multiple tasks. These works commonly trained the ViT [69] model or its variants with large quantities of parameters on huge amounts of data. The parameter-efficient transfer learning (PETL) [21], a.k.a. parameter-efficient fine-tuning (PEFT), first proposed in natural language processing, is an effective way to adapt foundation models to various downstream tasks, which can reduce trainable parameters and avoid catastrophic forgetting. The main PETL methods fall broadly into three categories: adding task-specific adapters [21], prompt tuning [22], and Low-Rank Adaptation (LoRA) [70]. Our previous works [13, 24] followed the first to adapt the pre-trained foundation models to the VPR task, while they are efficient only in terms of parameters, not in training time and GPU memory usage. Fortunately, there are some recent works [23, 29] that have preliminarily attempted to explore memory-efficient transfer learning methods. Inspired by them, we design a memory-efficient adaptation architecture tailored for the VPR task. Concretely, we use the tunable adapters to refine the intermediate features from the frozen backbone instead of inserting them into it to avoid backpropagation through the backbone, and upgrade vanilla adapters to MultiConv adapters to introduce the local priors and facilitate feature interactions along the spatial axes. As a result, we can achieve parameter-, time-, and memory-efficient adaptation and get highly robust VPR models.
@@ III-A Preliminary
where $x_{l-1}$ and $x_{l}$ are the output of the (l1)(l-1)-th and $l$-th transformer block.
For the feature map output by CNN models, a common practice of the VPR method is to use NetVLAD [6] or GeM pooling [61] to aggregate it into a global feature to conduct nearest neighbor search (i.e., global retrieval). For the ViT model, the output consists of one class token and $N$ patch tokens, where the class token can be directly used as the global feature to represent places [2, 59]. Meanwhile, $N$ patch tokens can also be reshaped as a feature map (similar to CNN). In this work, instead of using class token, we transform the feature map into the global feature using GeM pooling.
@@ III-B Memory-Efficient MultiConv Adaptation
However, due to the adapters inserted in the transformer block, this will lead to significant GPU memory consumption during training. Here, we conduct a brief equation derivation and theoretical analysis. Consider a transformer network with $L$ blocks, each containing the frozen parameters $\theta_{i}$ (frozen modules) and trainable parameters $\alpha_{i}$ (trainable modules), and producing the output $x_{i}$, which depends on both parameters $\theta_{i},\alpha_{i}$. What we need to do is to minimize a loss function $\mathcal{L}$ based on stochastic gradient descent. To be specific, the gradient for the trainable parameters $\alpha_{i}$ is calculated using the chain rule during backpropagation, formulated as follows:
@@
To this end, inspired by previous works [23, 29], we introduce a distinctive adaptation method. Instead of inserting the adapters into the backbone, we use a parallel network that does not require gradient backpropagation through the backbone during training. It is a simple yet effective architecture first proposed in previous work [23], which directly takes the intermediate features from the frozen backbone as input and progressively refines them (using lightweight transformer in [23] but MultiConv adapters in ours) for better representation. Concretely, we adapt the ViT-based foundation model DINOv2 [19] as the backbone, which is composed of a patch embedding layer and $L$ transformer blocks, and thereby $L+1$ intermediate features, $x_{0},x_{1},x_{2}...,x_{L}$, each consisting of $N+1$ tokens with a dimension of $D$, so $x_{i}\in\mathcal{R}^{(N+1)\times D}$ as mentioned in III-A. To refine them progressively, we construct the parallel adapters alongside the backbone. Meanwhile, we add residual connections to each adapter for better performance. We represent the output of the $l$-th parallel adapter as $y_{l}$. The whole computation can be formulated as:
@@
In this work, we introduce the multi-scale convolution (MultiConv) adapter, which is first proposed in our previous CricaVPR work [24] but is used by inserting it into the transformer block (similar to SelaVPR but different from SelaVPR++). Notably, in CricaVPR, the MultiConv adapter is used to add multi-scale local prior knowledge to the model for enhancing the performance in the VPR task. However, in SelaVPR++, it not only introduces multi-scale local priors, but more importantly, it breaks the limitation that the vanilla adapter cannot model the interactions between different patch tokens. In the MultiConv adapter, we add a MultiConv module (with a skip connection) between the activation layer and the up-projection layer, as shown in Fig. 3 (c). This module contains three parallel convolution paths of different scales ($\times1$, $\times3$, $\times5$), which is inspired by the inception module in GoogLeNet [73]. The $\times1$ convolution is also applied before the $\times3$ and $\times5$ convolutions to reduce the number of channels. This design and the bottleneck structure make our MultiConv adapter still lightweight. The outputs of the three convolution paths are subsequently concatenated to achieve integration. The MultiConv adapter only processes patch tokens, which are reshaped into a $W\times H\times D$-dim (from $N\times D$) feature map to restore the spatial position before being input into the adapter. In this way, the spatially adjacent patch tokens can be modeled interactions via $\times3$ and $\times5$ convolutions. Moreover, since the MultiConv adapter has introduced multi-scale local prior knowledge to the model, we no longer need to purposefully design a separate local adaptation module to process local information for VPR as in our previous SelaVPR.
It is worth noting that both the work [29] and our work follow the memory-efficient adaptation paradigm proposed in previous work [23]. Apart from the slight differences in overall pipeline (e.g., the use of residual connections parallel to the added trainable modules, and the order in which features from backbone are refined), the key difference between the three lies in the design of the trainable modules: the work [23] uses lightweight transformer blocks; the work [29] uses token-channel alternating adapters; we use MultiConv adapters.
@@ III-D Similarity-Constrained Deep Hashing
Hashing has gained widespread attention in approximate nearest neighbor search for large-scale image retrieval due to its computational efficiency and retrieval quality by encoding data into binary vectors and performing search with Hamming distance. Deep hashing, in particular, has further enhanced performance by combining deep representation learning with hash coding. In this work, we attempt to adopt deep hashing to provide compact binary place representation, which is achieved by using a sign function $sgn(\cdot)$ to quantize floating-point features $\boldsymbol{f}_{i}$} into binary hash codes $\boldsymbol{b}_{i}$}. Formally,
@@
In contrast to [74], which applies the sigmoid transformation on inner products to reduce the sensitivity to large values in the features, our method directly preserves the cosine similarity of pair-wise features and can be simplified into a more concise form (without sigmoid). Due to the advanced backbone and metric loss to produce well-behaved features, as well as L2-norm on features in our work, we no longer consider the issue of harmful large values in features as in [74].
After obtaining $\mathcal{L}_{Q}(\boldsymbol{f},\boldsymbol{b})$, to train a deep hashing model, we can use the total loss $\mathcal{L}$ that combines the metric loss $\mathcal{L}_{M}(\boldsymbol{f})$ and similarity-constrained loss $\mathcal{L}_{Q}(\boldsymbol{f},\boldsymbol{b})$ by a weight $\lambda$ as:
@@
Another solution to address the gradient issue of quantization operation is straight-through estimation (STE) [28], which was proposed in the image compression area and addressed the gradient issue of rounding quantization function. Since the sign function in deep hashing is also a quantization function (which can be seen as a special form of rounding function), we employ it to perform the backpropagation of deep hashing in the VPR task. Specifically, we use the original sign function $sgn(\cdot)$ during forward propagation to obtain the binary hash codes. In back propagation, the derivative of binary hashing operation is replaced with the derivative of a smooth approximation $g$, which can be formulated as: